Ricardo throws a stone off a bridge into a river below. The stone's height above the water in meters, y, is dependent on time in seconds after throwing, x, and can be modeled with the function y=−5(x−8)(x+4).
How many seconds after being thrown will the stone hit the water? seconds

What is the maximum height that the stone will reach? meters

How many seconds after being thrown will the stone reach its maximum height? seconds

What is the height of the stone at the time it is thrown? meters

Answer by stanbon(75887)

Ricardo throws a stone off a bridge into a river below. The stone's height above the water in meters, y, is dependent on time in seconds after throwing, x, and can be modeled with the function
y=−5(x−8)(x+4)
------
y = -5(x^2 -4x -32) = -5x^2 + 20x + 160
---------------------------------
How many seconds after being thrown will the stone hit the water? seconds
Solve -5(x-8)(x+4) = 0
Ans: x = 8 seconds
---------------------------
What is the maximum height that the stone will reach? meters
Vertex occurs at x = -b/(2a) = -20/(-10) = 2 seconds
Max height = f(2) = -5(-6)(6) = 180 meters
---------------------
How many seconds after being thrown will the stone reach its maximum height? seconds
Ans: 2 seconds (from vertex work above)
-----------------

What is the height of the stone at the time it is thrown?
Ans: f(0) = 160 meters
============
Cheers,
Stan H.
---------